# Using negative controls to identify causal effects with invalid instrumental variables

### v4

@@ A.2 Results for hypothesis testing
$\forall{z}\in\{0,1\},c\in\mathcal{C}_{0}\cap\mathcal{C}_{1}$, then it follows that
$E\{Y-t(Z,\mathbf{c})|Z=1,S=1,\mathbf{c}\}=E\{Y-t(Z,\mathbf{c})|Z=0,S=1,\mathbf{c}\}.$ (15)
@@ A.3 Negative control outcomes - identification
$E\{Y^{0}|A=1,Z,\mathbf{C}\}-E\{Y^{0}|A=0,Z,\mathbf{C}\}=\gamma(\mathbf{C})$ (16)
@@
$E\{Y^{1}-Y^{0}|A=1,Z=z,\mathbf{C}=\mathbf{c}\}$
$=E\{Y|A=1,z,\mathbf{c}\}-E\{Y|A=0,z,\mathbf{c}\}$
$\quad+\frac{E\{Y|A=0,Z=1,\mathbf{c}\}-E\{Y|A=0,Z=0,\mathbf{c}\}}{E(A|Z=1,\mathbf{c})-E(A|Z=0,\mathbf{c})}-\frac{E\{W|Z=1,\mathbf{c}\}-E\{W|Z=0,\mathbf{c}\}}{E(A|Z=1,\mathbf{c})-E(A|Z=0,\mathbf{c})}.$
@@ A.4 Assumptions for a general instrumental variable model
Consistency: $Y=Y^{\mathbf{A}}$ with probability 1 in those with $S=1$.
@@
with probability 1.
@@ A.5 Estimation under NSM
$q(\mathbf{a},\mathbf{Z,C})=E(Y^{\mathbf{0}}|\mathbf{A=a,Z},S=1,\mathbf{C})-E(Y^{\mathbf{0}}|\mathbf{A=0,Z},S=1,\mathbf{C})$
satisfies the restriction $q(\mathbf{a,Z,C})=q(\mathbf{a,C})$.
@@
where
$b_{1}(\mathbf{C})\equiv E(Y^{\mathbf{0}}|\mathbf{Z=0},S=1,\mathbf{C})-E(Y^{\mathbf{0}}|\mathbf{Z=0},S=0,\mathbf{C})$
$b_{0}(\mathbf{C})\equiv E(Y^{0}|\mathbf{Z=0},S=0,\mathbf{C}).$
@@
$\Gamma=\left\{\kappa(\mathbf{A,Z,C}):E\{\kappa(\mathbf{A,Z,C})|\mathbf{Z},S=1,\mathbf{C}\}=E\{\kappa(\mathbf{A,Z,C})|\mathbf{A},S=1,\mathbf{C}\}=0\right\}.$
@@
Under the NSM assumption, we will utilise an additional parametric model $q(\mathbf{A,C};\omega^{\dagger})$ for the selection bias function $q(\mathbf{A,C})$, where $q(\mathbf{A,C};\omega^{\dagger})$ is a known function smooth in a finite dimensional parameter $\omega^{\dagger}$. A smooth model $f(\mathbf{A}|\mathbf{Z},S=1,\mathbf{C};\pi^{\dagger})$ is postulated for $f(\mathbf{A,Z}|S=1,\mathbf{C})$; the event that model $f(A|\mathbf{Z},S,\mathbf{C};\pi^{\dagger})$ is correctly specified is denoted by $\mathcal{M}_{a}$. We will also consider alternative versions of the previous restrictions $\mathcal{M}_{1q}$ and $\mathcal{M}_{3q}$ on the data:
@@
  • $\mathcal{M}_{3q}$: $f(S|\mathbf{Z}=0,\mathbf{C})=f(S|\mathbf{Z}=0,\mathbf{C};\alpha^{\dagger})$, $OR(\mathbf{Z},S,\mathbf{C})=OR(\mathbf{Z},S,\mathbf{C};\rho^{\dagger})$, $b_{1}(\mathbf{C})=b_{1}(\mathbf{C};\theta^{\dagger}_{1})$, and $q(\mathbf{A,C})=q(\mathbf{A,C};\omega^{\dagger})$.
@@ A.6 Algorithm for constructing cross-fit de-biased machine learning estimators
    $0=\sum_{i\in I_{k}}\phi(\mathbf{Z}_{i},S_{i},\mathbf{C}_{i};\hat{\eta}^{c}_{k})\epsilon_{i}^{*}(\psi^{\dagger},\hat{\eta}_{k})$
@@
and $\epsilon^{*^{\prime}}(\psi,\eta)=\partial\epsilon^{*}(\psi,\eta)/\partial\psi$ is a row vector that is the length of the the dimension of $\psi$.
As in other semiparametric regression problems, a complication is how to perform machine-learning based estimation of the nuisance parameters, given that they are not all conditional mean functions that can be estimated directly; see Section 4.2 of Chernozhukov et al., (2018) for further discussion. One option is to restrict to estimation methods that can more easily respect the structure of the semiparametric model (generalised additive models, Lasso, deep neural networks etc); the other is to adopt the ‘localised de-biased machine learning’ procedure described in Kallus et al., (2019), where e.g. an initial (biased) estimator of $\psi^{\dagger}$ could be used to construct estimators of other nuisance parameters.
@@ A.7 Estimation strategy for $\psi^{\dagger}$ based on parametric working models
    $0=\sum^{n}_{i=1}$ $\left[e_{1}(\mathbf{Z}_{i},\mathbf{C}_{i})-\frac{E\{e_{1}(\mathbf{Z}_{i},\mathbf{C}_{i})f(S_{i}=1|\mathbf{Z}_{i},\mathbf{C}_{i};\hat{\alpha},\rho^{\dagger})f(S_{i}=0|\mathbf{Z}_{i},\mathbf{C}_{i};\hat{\alpha},\rho^{\dagger})|\mathbf{C}_{i};\hat{\alpha},\hat{\tau},\rho^{\dagger}\}}{E\{f(S_{i}=1|\mathbf{Z}_{i},\mathbf{C}_{i};\hat{\alpha},\rho^{\dagger})f(S_{i}=0|\mathbf{Z}_{i},\mathbf{C}_{i};\hat{\alpha},\rho^{\dagger})|\mathbf{C}_{i};\hat{\alpha},\hat{\tau},\rho^{\dagger}\}}\right]$
@@
    $0=\sum^{n}_{i=1}$ $(1-S_{i})\begin{pmatrix}e_{2}(\mathbf{Z}_{i},\mathbf{C}_{i})\\ e_{3}(\mathbf{C}_{i})\end{pmatrix}\epsilon^{*}_{i}(\psi^{\dagger},\nu^{\dagger},\theta^{\dagger})$
@@
    for $\theta_{1}^{\dagger}$ and $\psi^{\dagger}$, where $e_{4}(\mathbf{C}_{i})$ and $e_{5}(\mathbf{Z}_{i},\mathbf{C}_{i})$ are of the same dimension as $\theta_{1}^{\dagger}$ and $\psi^{\dagger}$ respectively.
@@
    $0=\sum^{n}_{i=1}\phi(\mathbf{Z}_{i},S_{i},\mathbf{C}_{i};\hat{\tau},\hat{\alpha},\hat{\rho})\epsilon^{*}(\psi^{\dagger},\hat{\nu},\hat{\theta})$
@@
for $\omega^{\dagger}$. Then $\hat{\psi}_{MR-NSM}$ is obtained by solving the equations
$0=\sum^{n}_{i=1}$ $\begin{pmatrix}\phi(\mathbf{Z}_{i},S_{i},\mathbf{C}_{i};\hat{\tau},\hat{\alpha},\hat{\rho})\epsilon_{i}^{*}(\psi^{\dagger},\hat{\nu},\hat{\theta})\\ S\kappa(\mathbf{A,Z,C};\hat{\pi},\hat{\tau})\epsilon_{i}(\psi^{\dagger},\hat{\nu},\hat{\theta},\hat{\omega},\hat{\pi})\end{pmatrix}$
@@ A.8 Local semiparametric efficiency
$\mu(\mathbf{Z},S,\mathbf{C})=E\left\{\frac{\partial\beta(\mathbf{A,Z,C};\psi^{\dagger})S}{\partial\psi}\bigg{|}_{\psi=\psi^{\dagger}}\bigg{|}\mathbf{Z},S,\mathbf{C}\right\},$
$P_{\sigma}(\mathbf{Z,C})=\frac{E\{S\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}{E\{\sigma^{-2}(\mathbf{Z},S,\mathbf{C})|\mathbf{Z,C}\}}$
and $\delta\mu(\mathbf{Z,C})=\mu(\mathbf{Z},1,\mathbf{C})-\mu(\mathbf{Z},0,\mathbf{C})$, where $\sigma^{2}(\mathbf{Z},S,\mathbf{C})=E\{\epsilon^{*2}(\psi^{\dagger})|\mathbf{Z},S,\mathbf{C}\}$.
@@
Under the semiparametric model $\mathcal{M}_{NEM}$ and with admissible independence densities
@@
the optimal index function $r^{opt}_{0}(\mathbf{Z},S,\mathbf{C})$ in (8) is
$r_{0}^{opt}(\mathbf{Z},S,\mathbf{C})=\left\{\delta\mu(\mathbf{Z,C})-\frac{E\left\{\delta\mu(\mathbf{Z},\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}\right\}2(-1)^{1-S}P_{\sigma}(\mathbf{Z,C}).$
@@
A proof is given in B.8 of the Appendix. We emphasise that the choice of admissible independence densities is not an assumption on the data generating mechanism. Rather, it allows us to simplify the complex expression for the efficient score given in the Appendix into an equation of the form (8). Similar to the standard instrumental variable set-up, efficient estimation generally requires modelling the conditional mean of $\mathbf{A}$ given $\mathbf{Z}$ and $\mathbf{C}$ in those with $S=1$. However, even if this model is misspecified, the resulting estimating of $\psi^{\dagger}$ will remain consistent (although it will no longer generally be efficient). In order to further simplify the above result, we revisit Example 1.
@@
$m_{{opt}}(\mathbf{C})=E\left[\left\{\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\right\}^{2}\sigma^{2}(\mathbf{Z},S,\mathbf{C})\bigg{|}\mathbf{C}\right]^{-1}\{\mu(Z=1,S=1,\mathbf{C})-\mu(Z=0,S=1,\mathbf{C})\}.$
If $\sigma^{2}(\mathbf{Z},S,\mathbf{C})=\sigma^{2}$, then $m_{{opt}}(\mathbf{C})$ reduces to $w_{0}(\mathbf{C})^{-1}\sigma^{-2}\{\mu(Z=1,S=1,\mathbf{C})-\mu(Z=0,S=1,\mathbf{C})\}$, where
$w_{0}(\mathbf{C})=\left\{\frac{1}{f(Z=1,S=1|\mathbf{C})}+\frac{1}{f(Z=1,S=0|\mathbf{C})}+\frac{1}{f(Z=0,S=1|\mathbf{C})}+\frac{1}{f(Z=0,S=0|\mathbf{C})}\right\}.$
@@ A.9 Negative control outcomes - estimation
$q(\mathbf{a,Z,C})=E(Y^{\mathbf{0}}|\mathbf{A=a},\mathbf{Z},\mathbf{C})-E(Y^{\mathbf{0}}|\mathbf{A=0},\mathbf{Z},\mathbf{C})$
$t(\mathbf{z,C})=E(Y^{\mathbf{0}}|\mathbf{Z=z},\mathbf{C})-E(Y^{\mathbf{0}}|\mathbf{Z=0},\mathbf{C})$
$b^{*}_{1}(\mathbf{C})=E(Y^{\mathbf{0}}|\mathbf{Z=0},\mathbf{C})$
$b_{0}(\mathbf{C})=E(W^{\mathbf{0}}|\mathbf{Z=0},\mathbf{C}).$

### v6

@@ A.2 Results for hypothesis testing
$\forall{z}\in\{0,1\},\mathbf{c}\in\mathcal{C}_{0}\cap\mathcal{C}_{1}$, then it follows that
$E\{Y-t(Z,\mathbf{c})|Z=1,S=1,\mathbf{C}=\mathbf{c}\}=E\{Y-t(Z,\mathbf{c})|Z=0,S=1,\mathbf{C}=\mathbf{c}\}.$ (13)
@@ A.3 Negative control outcomes - identification
$E(Y^{0}|A=1,Z,\mathbf{C})-E(Y^{0}|A=0,Z,\mathbf{C})=\gamma(\mathbf{C})$ (14)
@@
$E(Y^{1}-Y^{0}|A=1,Z=z,\mathbf{C}=\mathbf{c})$
$=E(Y|A=1,z,\mathbf{c})-E(Y|A=0,z,\mathbf{c})$
+$\quad+\left\{\frac{E(Y|A=0,Z=1,\mathbf{c})-E(Y|A=0,Z=0,\mathbf{c})}{E(A|Z=1,\mathbf{c})-E(A|Z=0,\mathbf{c})}\right\}-\left\{\frac{E(W|Z=1,\mathbf{c})-E(W|Z=0,\mathbf{c})}{E(A|Z=1,\mathbf{c})-E(A|Z=0,\mathbf{c})}\right\}.$
@@ A.4 Assumptions for a general instrumental variable model
Consistency: $Y=Y^{\mathbf{A}}$ with probability 1 in individuals with $S=1$.
@@
with probability 1, for all $\mathbf{z}$ in the relevant support of $\mathbf{Z}$.
@@ A.5 Estimation under NSM
$q(\mathbf{a},\mathbf{Z,C})\equiv E(Y^{\mathbf{0}}|\mathbf{A=a,Z},S=1,\mathbf{C})-E(Y^{\mathbf{0}}|\mathbf{A=0,Z},S=1,\mathbf{C})$
satisfies the restriction $q(\mathbf{a,Z,C})=q(\mathbf{a,C})$ with probability 1.
@@
where we redefine
$b_{1}(\mathbf{C})\equiv$ $E\left(Y-\beta(\mathbf{A,Z,C})S-[q(\mathbf{A,Z,C})-E\{q(\mathbf{A,Z,C})|\mathbf{Z},S=1,\mathbf{C}\}]S|\mathbf{Z=0},S=1,\mathbf{C}\right)$
$-E(Y|\mathbf{Z=0},S=0,\mathbf{C})$
$b_{0}(\mathbf{C})\equiv$ $E(Y|\mathbf{Z=0},S=0,\mathbf{C}).$
@@
$\Gamma\equiv\left\{\kappa(\mathbf{A,Z,C}):E\{\kappa(\mathbf{A,Z,C})|\mathbf{Z},S=1,\mathbf{C}\}=E\{\kappa(\mathbf{A,Z,C})|\mathbf{A},S=1,\mathbf{C}\}=0\right\}.$
@@
Under the NSM assumption, we will utilise an additional parametric model $q(\mathbf{A,C};\omega^{\dagger})$ for the selection bias function $q(\mathbf{A,C})$, where $q(\mathbf{A,C};\omega^{\dagger})$ is a known function smooth in a finite dimensional parameter $\omega^{\dagger}$. A smooth model $f(\mathbf{A}|\mathbf{Z},S=1,\mathbf{C};\pi^{\dagger})$ is postulated for $f(\mathbf{A}|\mathbf{Z},S=1,\mathbf{C})$; the event that model $f(\mathbf{A}|\mathbf{Z},S=1,\mathbf{C};\pi^{\dagger})$ is correctly specified is denoted by $\mathcal{M}_{a}$. We will also consider alternative versions of the previous restrictions $\mathcal{M}_{1q}$ and $\mathcal{M}_{3q}$ on the data:
@@
  • $\mathcal{M}_{3q}$: $f(S|\mathbf{Z}=\mathbf{0},\mathbf{C})=f(S|\mathbf{Z}=\mathbf{0},\mathbf{C};\alpha^{\dagger})$, $OR(\mathbf{Z},S,\mathbf{C})=OR(\mathbf{Z},S,\mathbf{C};\rho^{\dagger})$, $b_{1}(\mathbf{C})=b_{1}(\mathbf{C};\theta^{\dagger}_{1})$, and $q(\mathbf{A,C})=q(\mathbf{A,C};\omega^{\dagger})$.
@@ A.6 Algorithm for constructing cross-fit de-biased machine learning estimators
    $0=\sum_{i\in I_{k}}\phi(\mathbf{Z}_{i},S_{i},\mathbf{C}_{i};\hat{\eta}^{c}_{k})\epsilon_{i}^{*}(\psi^{\dagger},\hat{\eta}^{c}_{k})$
@@
and $\epsilon^{*^{\prime}}(\psi,\eta)=\partial\epsilon^{*}(\psi,\eta)/\partial\psi$ is a row vector that is the length of the dimension of $\psi$.
As in other semiparametric regression problems, a complication is how to perform machine learning-based estimation of the nuisance parameters, given that they are not all conditional mean functions that can be estimated directly; see Section 4.2 of Chernozhukov et al., (2018) for further discussion. One option is to restrict to estimation methods that can more easily respect the structure of the semiparametric model (generalised additive models, Lasso, deep neural networks etc); the other is to adopt the ‘localised de-biased machine learning’ procedure described in Kallus et al., (2019), where e.g. an initial (biased) estimator of $\psi^{\dagger}$ could be used to construct estimators of other nuisance parameters.
@@ A.7 Estimation strategy for $\psi^{\dagger}$ based on parametric working models
    $0=\sum^{n}_{i=1}$ $\left[e_{1}(\mathbf{Z}_{i},\mathbf{C}_{i})-\frac{E\{e_{1}(\mathbf{Z}_{i},\mathbf{C}_{i})f(S_{i}=1|\mathbf{Z}_{i},\mathbf{C}_{i};\hat{\alpha},\rho^{\dagger})f(S_{i}=0|\mathbf{Z}_{i},\mathbf{C}_{i};\hat{\alpha},\rho^{\dagger})|\mathbf{C}_{i};\hat{\tau},\hat{\alpha},\rho^{\dagger}\}}{E\{f(S_{i}=1|\mathbf{Z}_{i},\mathbf{C}_{i};\hat{\alpha},\rho^{\dagger})f(S_{i}=0|\mathbf{Z}_{i},\mathbf{C}_{i};\hat{\alpha},\rho^{\dagger})|\mathbf{C}_{i};\hat{\tau},\hat{\alpha},\rho^{\dagger}\}}\right]$
@@
    $0=\sum^{n}_{i=1}$ $(1-S_{i})\begin{pmatrix}e_{2}(\mathbf{Z}_{i},\mathbf{C}_{i})\\ e_{3}(\mathbf{C}_{i})\end{pmatrix}\epsilon^{*}_{i}(\psi^{\dagger},\nu^{\dagger},\theta^{\dagger}_{0})$
@@
    for $\theta_{1}^{\dagger}$ and $\psi^{\dagger}$, where $e_{4}(\mathbf{C})$ and $e_{5}(\mathbf{Z},\mathbf{C})$ are of the same dimension as $\theta_{1}^{\dagger}$ and $\psi^{\dagger}$ respectively.
@@
    $0=\sum^{n}_{i=1}\phi(\mathbf{Z}_{i},S_{i},\mathbf{C}_{i};\hat{\tau},\hat{\alpha},\hat{\rho})\epsilon_{i}^{*}(\psi^{\dagger},\hat{\nu},\hat{\theta})$
@@
for $\omega^{\dagger}$, where $e_{6}(A,\mathbf{C})$ is of the same dimension as $\omega^{\dagger}$. Then $\hat{\psi}_{MR-NSM}$ is obtained by solving the equations
$0=\sum^{n}_{i=1}$ $\begin{pmatrix}\phi(\mathbf{Z}_{i},S_{i},\mathbf{C}_{i};\hat{\tau},\hat{\alpha},\hat{\rho})\epsilon_{i}^{*}(\psi^{\dagger},\hat{\nu},\hat{\theta})\\ S\kappa(\mathbf{A}_{i},\mathbf{Z}_{i},\mathbf{C}_{i};\hat{\pi},\hat{\tau})\epsilon_{i}(\psi^{\dagger},\hat{\nu},\hat{\theta},\hat{\omega},\hat{\pi})\end{pmatrix}$
@@ A.8 Local semiparametric efficiency
$\mu(\mathbf{Z},S,\mathbf{C})\equiv E\left\{\frac{\partial\beta(\mathbf{A,Z,C};\psi^{\dagger})S}{\partial\psi}\bigg{|}_{\psi=\psi^{\dagger}}\bigg{|}\mathbf{Z},S,\mathbf{C}\right\},$
$P_{\sigma}(\mathbf{Z,C})\equiv\frac{f(S=1|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},1,\mathbf{C})f(S=0|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},0,\mathbf{C})}{f(S=1|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},1,\mathbf{C})+f(S=0|\mathbf{Z,C})\sigma^{-2}(\mathbf{Z},0,\mathbf{C})},$
where $\sigma^{2}(\mathbf{Z},S,\mathbf{C})\equiv E\{\epsilon^{*2}(\psi^{\dagger})|\mathbf{Z},S,\mathbf{C}\}$.
@@
Under the semiparametric model $\mathcal{M}_{NEM}$ and with admissible independence density
@@
the optimal index function $r^{opt}_{0}(\mathbf{Z},S,\mathbf{C})$ in
is
$r_{0}^{opt}(\mathbf{Z},S,\mathbf{C})=\left\{\mu(\mathbf{Z},1,\mathbf{C})-\frac{E\left\{\mu(\mathbf{Z},1,\mathbf{C})P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}{E\left\{P_{\sigma}(\mathbf{Z,C})|\mathbf{C}\right\}}\right\}2(-1)^{1-S}P_{\sigma}(\mathbf{Z,C}).$
@@
A proof is given in B.8. We emphasise that the choice of admissible independence densities is not an assumption on the data generating mechanism. Rather, it allows us to simplify the complex general expression for the efficient score into an equation of the form (15). Similar to the standard instrumental variable set-up, efficient estimation generally requires modelling the conditional mean of $\mathbf{A}$ given $\mathbf{Z}$ and $\mathbf{C}$ in those with $S=1$. However, even if this model is misspecified, the resulting estimating of $\psi^{\dagger}$ will remain consistent (although it will no longer generally be efficient). In order to further simplify the above result, we revisit Example 1.
@@
$m_{{opt}}(\mathbf{C})=E\left[\left\{\frac{(-1)^{Z+S}}{f(Z,S|\mathbf{C})}\right\}^{2}\sigma^{2}(Z,S,\mathbf{C})\bigg{|}\mathbf{C}\right]^{-1}\{\mu(Z=1,S=1,\mathbf{C})-\mu(Z=0,S=1,\mathbf{C})\}.$
If $\sigma^{2}(Z,S,\mathbf{C})=\sigma^{2}$, then $m_{{opt}}(\mathbf{C})$ reduces to $w_{0}(\mathbf{C})^{-1}\sigma^{-2}\{\mu(Z=1,S=1,\mathbf{C})-\mu(Z=0,S=1,\mathbf{C})\}$, where
$w_{0}(\mathbf{C})=\frac{1}{f(Z=1,S=1|\mathbf{C})}+\frac{1}{f(Z=1,S=0|\mathbf{C})}+\frac{1}{f(Z=0,S=1|\mathbf{C})}+\frac{1}{f(Z=0,S=0|\mathbf{C})}.$
@@ A.9 Negative control outcomes - estimation
$q(\mathbf{a,Z,C})\equiv$ $E(Y^{\mathbf{0}}|\mathbf{A=a},\mathbf{Z},\mathbf{C})-E(Y^{\mathbf{0}}|\mathbf{A=0},\mathbf{Z},\mathbf{C})$
$t(\mathbf{z,C})\equiv$ $E(Y^{\mathbf{0}}|\mathbf{Z=z},\mathbf{C})-E(Y^{\mathbf{0}}|\mathbf{Z=0},\mathbf{C})$
$b^{*}_{1}(\mathbf{C})\equiv$ $E(Y^{\mathbf{0}}|\mathbf{Z=0},\mathbf{C})$
$b_{0}(\mathbf{C})\equiv$ $E(W^{\mathbf{0}}|\mathbf{Z=0},\mathbf{C}).$